OpenStudy (anonymous):

3(a+b)-2(a+b) What is it simplified and how?

OpenStudy (anonymous):

=3a+3b-2a-2b=a+b

OpenStudy (bahrom7893):

(a+b)(3-2) = (a+b)

OpenStudy (anonymous):

the above answers r correct. you apply the numbers outside the parentheses to each figure inside them. then u just simplify like figures